OpenStudy (anonymous):

(medal given) For each function find the natural domain a.f(x)=6x^2-x-2 b.f(x)=square root x-3 c.f(x)=1/square root x-3 d.f(x)=1/x +1

OpenStudy (anonymous):

wanna know how to solve it or he answers?

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

A) \[-\infty,\infty\]

OpenStudy (anonymous):

Those will be in parentheses like (-infinty, infinty)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

B. (3, ∞)

OpenStudy (anonymous):

C.) same as B D) (-∞, 1) (1,∞)

OpenStudy (anonymous):

Ignore B It would be [3, ∞) for both B and C, the type of parentheses used is important

OpenStudy (anonymous):

\[for B ,\it is [3,\infty) .for C \it is \left( 3, \infty\right) \]

OpenStudy (anonymous):

\[for D ,\left( -\infty,-1 \right)\cup \left( -1,\infty \right) \] or all real numbers except -1

OpenStudy (anonymous):

@surjithayer is right about those

OpenStudy (anonymous):

1.denominator should not be zero. 2.terms under radical sign should be >=0